How do you find all Real Solutions for this equation?
|3x+5| = 1-x

Answer by jim_thompson5910(35256)

|3x+5| = 1-x

3x+5 = (1-x) or 3x+5 = -(1-x)

3x+5 = 1-x or 3x+5 = -1+x

3x = 1-x-5 or 3x = -1+x-5

3x+x = 1-5 or 3x-x = -1-5

4x = -4 or 2x = -6

x = -4/4 or x = -6/2

x = -1 or x = -3

The two solutions are x = -1 or x = -3


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim